OpenStudy (anonymous):

the volume of an equilateral triangular prism must be 420cm^3. Find the dimensions which will use the least construction materials. (minimize surface area) hawlp ;c

OpenStudy (bahrom7893):

http://0.tqn.com/d/math/1/0/J/F/Triangularprismr.gif it looks like that, only b=s

OpenStudy (anonymous):

thank you kind sir, you have saved me from calculus hell!

OpenStudy (bahrom7893):

V = (1/2)s*h = 420, so s*h = 840

OpenStudy (bahrom7893):

A = s*h+2l*s + l*s, again b=s, I'm just using the formulas from that pic, but that pic has an isosceles triangular prism, and u want equilateral

OpenStudy (anonymous):

yes i need equilateral. but how exactly do i fidn the dimensions if i wasnt given any other information besides the volume?

OpenStudy (bahrom7893):

@myininaya i've gotta go now, can u finish helping her? thanks Myin